(48) Business and finance. When x number of units are sold, the price of each unit (in dollars) is given by . Find the unit price when the following quantities are sold: 2,7,9,11.
(64) Business and finance. The revenue for a sandwich shop is directly proportional to its advertising budget. When the owner spent $2000 a month on advertising, the revenue was $120,000. If the revenue is now $180,000, how much is the owner spending on advertising?


Answer by stanbon(48554)

The price formula did not get posted.
--------------
The revenue for a sandwich shop is directly proportional to its advertising budget. When the owner spent $2000 a month on advertising, the revenue was $120,000. If the revenue is now $180,000, how much is the owner spending on advertising?
Formula: rev=k(ad bud)
120,000=k(2000)
k=60
Equation: rev=60*(ad bud)
180,000=60(ad bud)
advertising budet = $3000
Cheers,
Stan H.